OpenStudy (anonymous):

Find the particular solution of the differential equation )x^2dy)/(y^2-3dx)=(1/2y) satisfying the initial condition y(1)=(4)^(1/2)

OpenStudy (anonymous):

\[\frac{x^2 dy}{y^2-3 dx}=\frac{1}{2y}\] this one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(x^2 2y )dy-(y^2-3 dx)=0 \] Nx= 4 xy My= 2y so obviously not exact

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then

OpenStudy (anonymous):

?

OpenStudy (anonymous):

You may have to use special integration factor,

OpenStudy (anonymous):

like what?

OpenStudy (anonymous):

ok nevermind can u help with this one Solve the separable differential equation for u : du/dt =e^(6u+6t) use the following initial condition u(0)=10